# Properties of kaon-condensed phase in hyperon-mixed matter with three-baryon forces

### v1

@@ V.2 Role of three-body repulsion and attraction on the equation of state in SNM
The subtle difference between the $E$ (TNR) obtained from SJM2 and that from LP(1981) at $\rho_{\rm B}=\rho_{0}$ leads to substantial difference of repulsive energy at high densities $\rho_{\rm B}\gtrsim$ 0.40 fm-3 due to the nonlinear density-dependence of the $E$ (TNR) and $E$ (TNR) (LP) ($\propto\rho_{\rm B}^{2}$). There is also a rapid increase in $E$ (two-body) over the density $\rho_{\rm B}\sim$ 0.40 fm-3 as compared with the case of LP (1981). This is due to the fact that, in a meson-exchange picture like the MRMF model, attraction brought about by the $\sigma$-meson exchange is saturated at high densities, while repulsion by the $\omega$-meson exchange steadily increases. Such a relativistic effect further leads to stiffer EOS than the case of LP (1981).
For $\rho_{\rm B}\lesssim\rho_{0}$, the $E$ (two-body) in our case is less attractive than that of LP (1981) as seen in Fig. 2. The deficit attraction in the $E$ (two-body) is compensated with more attractive TNA than the case of LP (1981). For $\rho_{\rm B}\gtrsim\rho_{0}$, the $E$ (TNA) in both cases decrease in magnitude and become negligible for $\rho_{\rm B}\gtrsim$ 0.4 fm-3.
@@ V.3 Dependence of the EOS for SNM and PNM on the slope $L$
Next we consider dependence of the EOS for SNM on the slope $L$. In Fig. 3, the total energy per baryon as well as each energy contribution in SNM are shown as functions of $\rho_{\rm B}$ by the dotted, solid, and dashed lines for $L$=60, 65, and 70 MeV, respectively. The $E$ (TNR) is commonly taken from the SJM2 t2008 for the cases of $L$=(60, 65, 70) MeV and is represented by the single green solid line.
@@
At high densities beyond $\rho_{0}$, the attraction from the $\sigma$-exchange is saturated, so that the remaining repulsion from the $\omega$-exchange is more remarkable for larger $\omega$-mean field. The increase in repulsive energy from the $\omega$-meson exchange with the increase in $L$ from 60 MeV to 70 MeV, as shown by the arrows for the $E$ (two-body) in Fig. 3, plays a decisive role on stiffening the EOS in SNM.
To compare the $L$-dependence of the EOS between the PNM and SNM, the total energy per nucleon, $E$ (total), is shown in Fig. 4 and the pressure $p$ is shown in Fig. 5, as a function of $\rho_{\rm B}$ in the case of $L$=(60, 65, 70) MeV, for both PNM and SNM cases. One can see the EOS gets stiffer for larger $L$ over the densities $\rho_{\rm B}=(1-5)\rho_{0}$ for PNM as well as the case of SNM. Therefore, in our interaction model, not only the TNR but also the slope $L$ controls the stiffness of the EOS at high densities. This feature will be shown to be commonly applied to hadronic matter with the ($Y$+$K$) phase. (see Sec. VII. )
@@
Figure 5: The pressure in SNM (black curves) and in PNM (green curves) as a function of $\rho_{\rm B}$ for $L$=(60, 65, 70) MeV.
In danielewicz2002 , the allowable region in the pressure ($p$)-density ($\rho_{\rm B}$) plane for the EOS in SNM and PNM has been obtained from the analyses of transverse and elliptic flow data in nuclear collisions. Quantitatively, our result for $L\lesssim$ 65 MeV lie within the upper bound of the experimentally consistent region for $\rho_{\rm B}$ = (2$\sim$4) $\rho_{0}$, while those for $L\gtrsim$70 MeV slightly above it. For the latter case, some extra softening effect might be responsible in order to be consistent with the experimental implications.
In Fig. 4, we show the density-dependence of the symmetry energy, $S(\rho_{\rm B})$, defined by $S(\rho_{\rm B})$=$E$ (total)(PNM)--$E$ (total)(SNM), in the case of $L$=(60, 65, 70) MeV by the red solid lines. $S(\rho_{\rm B})$ increases monotonically with density for a given $L$. Since there is a cancellation of the $L$-dependent part of the total energy between the cases of PNM and SNM, the slope $L$-dependence of $S(\rho_{\rm B})$ is not remarkable even at high densities.
@@ VI Kaon properties in hyperon-mixed matter and onset of KC
Here we consider kaon properties in hyperon-mixed matter and obtain onset density of KC with our interaction model (ChL+MRMF+UTBR+TNA).
@@ VI.1 Meson-hyperon coupling constants for description of hyperon-mixed matter
For the description of hyperon-mixed matter, we set the values of the meson-hyperon coupling constants to obtain the hyperon-nucleon and hyperon-hyperon interactions in the MRMF.
@@
where $\langle m\rangle_{0}$ ($m$=$\sigma$, $\sigma^{\ast}$, $\omega$, $R_{0}$, $\phi_{0}$) are the meson mean fields at $\rho_{B}=\rho_{0}$ in SNM, and the last term, which is equal to the second term from the last on the r. h. s. of Eq. (79), comes from the energy density contribution from the UTBR. From recent theoretical and experimental analyses on hypernuclear experiments ghm2016 , the values of the $V_{Y}^{N}$ in Eq. (93) are set to be $V_{\Lambda}^{N}=-27$ MeV, $V_{\Sigma^{-}}^{N}$ = 23.5 MeV, and $V_{\Xi^{-}}^{N}$ = $-14$ MeV, and one obtains $g_{\sigma\Lambda}$, $g_{\sigma\Sigma^{-}}$, and $g_{\sigma\Xi^{-}}$. These coupling constants are obtained for each case of $L$ and listed in Table 2.
@@ VI.2 Onset of $\Lambda$-mixing and composition of matter below the onset density of KC
As we show below, the $\Lambda$ hyperons always appear at a lower density than KC for various cases of $L$ and $\Sigma_{Kn}$. Therefore first we consider the onset of the $\Lambda$-mixing and composition of matter prior to KC.
@@
Figure 6: (a)  The $\Lambda$ potential $V_{\Lambda}$ (green curves) and ($\mu_{n}-M_{\Lambda}$) (red curves) as functions of baryon number density $\rho_{\rm B}$ for $\Sigma_{Kn}$ = 300 MeV in the case of $L$=60 MeV (dotted lines), 65 MeV (solid lines), and 70 MeV (dashed lines). The filled triangle denotes the onset density of $\Lambda$-mixing, $\rho_{\rm B}(\Lambda)$, in $\beta$-equilibrated nuclear matter, and the filled circle denotes the onset density of KC, $\rho_{\rm B}^{c}(K^{-})$, in $\beta$-equilibrated hyperon-mixed matter. (b) The same as (a), but for $\Sigma_{Kn}$ = 400 MeV. The profiles for $\Sigma_{Kn}$ = 300 MeV and 400 MeV are the same before appearance of KC. See the text for details.
The $\rho_{\rm B}^{c}(\Lambda)$ in $\beta$-equilibrated neutron-star matter is given by the intersection of the curves for the $\mu_{n}-M_{\Lambda}$ and $V_{\Lambda}$, denoted as the filled triangle. The $\Lambda$ hyperons are mixed over the densities where $\mu_{n}-M_{\Lambda}>V_{\Lambda}$. Numerically $\rho_{\rm B}^{c}(\Lambda)$ is weakly dependent on $L$: $\rho_{\rm B}^{c}(\Lambda)$ = (0.47, 0.43, 0.40) fm-3 for $L$ = (60, 65, 70) MeV.
@@
Table 3: The onset densities at which hyperon-mixing starts and those of KC in the (ChL+MRMF+ UTBR+TNA) model for $\Sigma_{Kn}$ = 300 MeV and 400 MeV in case of $L$=60, 65, and 70 MeV. The $\rho_{\rm B}^{c}(\Lambda)$ is the onset density of $\Lambda$ hyperons in the normal neutron-star matter, $\rho_{\rm B}^{c}(\Xi^{-}\ {\rm in}\ \Lambda)$ the one of $\Xi^{-}$ hyperons in the $\Lambda$-mixed matter, $\rho_{\rm B}^{c}~{}(K^{-})$ the one of KC in the hyperon ($\Lambda$ and/or $\Xi^{-}$)-mixed matter, and $\rho_{\rm B}^{c}(\Xi^{-}\ {\rm in}\ K^{-}\Lambda)$ the one of the $\Xi^{-}$ hyperons in the KC phase in the $\Lambda$-mixed matter. For reference, the depth of the $K^{-}$ optical potential $U_{K}$, estimated at $\rho_{\rm B}$ = $\rho_{0}$ in the SNM, is listed for each case of $L$ and $\Sigma_{Kn}$.
$L$ $\Sigma_{Kn}$ $U_{K}$ $\rho_{\rm B}^{c}(\Lambda)$ $\rho_{\rm B}^{c}(\Xi^{-}~{}{\rm in}~{}\Lambda)$ $\rho_{\rm B}^{c}(K^{-})$ $\rho_{\rm B}^{c}(\Xi^{-}~{}{\rm in}~{}{K^{-}\Lambda})$
(MeV) (MeV) (MeV) (fm-3) (fm-3) (fm-3) (fm-3)
60 300 --111 0.466 -- 0.598 1.04
400 --132 0.466 -- 0.486 0.994
65 300 --111 0.425 0.568 0.653 --
400 --131 0.425 -- 0.503 0.900
70 300 --111 0.397 0.516 0.733 --
400 --131 0.397 (0.516) 0.523 0.790
@@ VI.3 Onset density of kaon condensation in hyperon-mixed matter
The lowest kaon energy $\omega_{K}(\rho_{B})$ decreases as a function of $\rho_{\rm B}$ due to the $K$-$B$ scalar and vector attraction. In the case of the $s$-wave kaon condensation, the $\omega_{K}(\rho_{B})$ intersects with the kaon chemical potential $\mu_{K}$, which is equal to the charge chemical potential $\mu$ (74). The onset density $\rho_{B}^{c}(K^{-})$ for the $s$-wave kaon condensation as a continuous phase transition is given by the condition mt92
@@
With the coefficients $d_{p}$ and $d_{n}$, which are determined so as to reproduce the empirical values of the on-shell $s$-wave $K$-$N$ scattering lengths (Appendix X.2), one can see from Eq. (99) that for on-shell $K^{-}$ mesons ($\omega_{K}=m_{K}$), the $s$-wave $K$-$N$ scalar attraction simulated by $\Sigma_{Kb}$ terms is canceled out by each corresponding range term. Also the $\Lambda^{\ast}$-pole, which lies $\sim$30 MeV below the $KN$ threshold, works repulsively to the on-shell $s$-wave $K^{-}$$p$ scattering length. The range terms and $\Lambda^{\ast}$-pole term thus have sizable contributions to the $K^{-}$ self-energy in free space. However, as baryon density increases such that $\rho_{\rm B}\gtrsim\rho_{0}$, the $K^{-}$ energy $\omega_{K}$ decreases due to the $s$-wave $K$-$N$ attractive interactions, getting $\omega_{K}/m_{K}\ll 1$. In such a case, the $\Sigma_{Kb}$ terms and $s$-wave $K$-$N$ vector interaction $-2X_{0}\omega_{K}$ become dominant over the range terms and the $\Lambda^{\ast}$-pole term [^2] In Ref. fmmt1996 , the same result has been obtained in the second-order perturbation with respect to the axial-vector current (the second-order effect) in the framework of current algebra and PCAC t88 . Our result on the lowest $K^{-}$ energy naturally agrees with those obtained by other works based on the chiral perturbation theory lbm95 ; tw1995 ., and the $K^{-}$ self-energy is well described without the range terms and $\Lambda^{\ast}$ term as
@@
which we use for actual numerical calculations throughout this paper. The term leading to the $s$-wave $K$-$B$ scalar interaction in the self-energy (100) is proportional to $\rho_{b}^{s}$, which has a complicated $\rho_{\rm B}$-dependence:
@@
A scale of the $s$-wave $K$-$N$ attraction is characterized by the $K^{-}$ optical potential $U_{K}$ in the SNM, which is defined in terms of the $K^{-}$ self-energy [(100)] as $U_{K}=\Pi_{K}(\omega_{K};\rho_{\rm B})/(2\omega_{K})|_{\rho_{\rm B}=\rho_{0}}$. In Table 3, the $U_{K}$ is listed for each case of $L$ and $\Sigma_{Kn}$. The value of $U_{K}$ has sensitive dependence on $\Sigma_{Kn}$, while it depends little on the slope $L$. Our deduced value of the depth $|U_{K}|$ (110 MeV--130 MeV) is larger than the theoretical values in the chiral unitary approach ro2000 and the recent optimal value of the real part of the $K^{-}$ optical potential depth $|V_{0}|$ = 80 MeV with the imaginary part $W_{0}$ = --40 MeV, which was deduced from the measured spectrum shape of the inclusive missing-mass spectrum of 12C ($K^{-},p$) reactions in the J-PARC E05 experiment ichikawa2020 . Nevertheless, our result is within the allowable range, considering that a much larger value of $|V_{0}|$ ($\gtrsim$ 150 MeV) is confronted with the difficulty of reproducing the spectrum ichikawa2020 .
@@
Going into details of (i), we show the density-dependence of the $X_{0}$ [Eq. (16)] and that of the “effective mass” $m_{K}^{\ast}$ of the $K^{-}$ meson [Eq. (65)] in Fig. 7. For reference, the particle fractions $\rho_{a}/\rho_{\rm B}$ and baryon scalar densities $\rho_{b}^{s}$ including the ($Y$+$K$) phase are also shown as functions of $\rho_{\rm B}$ in Fig. 8 and Fig. 9, respectively.
@@
In Refs. panda1995 ; carlson2000 the effect of $K$-$N$ and $N$-$N$ correlations was taken into account in the study of kaons in dense matter, and it was shown that the correlations tend to push up the $\omega_{K}$ over the relevant densities. Subsequently the effect of the short-range correlations was discussed in detail and was shown to be moderate wrw1997 ; ww1997 as compared to the preceding results panda1995 ; carlson2000 . Our result on the $U_{K}$ is similar to that of Refs. wrw1997 ; ww1997 with inclusion of the short-range correlations ($U_{K}\sim-120$MeV), whereas the $\omega_{K}$ in our model is slightly smaller than that of Refs. wrw1997 ; ww1997 .

### v2

@@ V.2 Role of three-body repulsion and attraction on the equation of state in SNM
The subtle difference between the $E$ (TNR) obtained from SJM2 and that from LP(1981) at $\rho_{\rm B}=\rho_{0}$ leads to substantial difference of repulsive energy at high densities $\rho_{\rm B}\gtrsim$ 0.40 fm-3 due to the nonlinear density-dependence of the $E$ (TNR) and $E$ (TNR) (LP) ($\propto\rho_{\rm B}^{2}$). There is also a rapid increase in $E$ (two-body) over the density $\rho_{\rm B}\approx$ 0.40 fm-3 as compared with the case of LP (1981). This is due to the fact that, in a meson-exchange picture like the MRMF model, attraction brought about by the $\sigma$-meson exchange is saturated at high densities, while repulsion by the $\omega$-meson exchange steadily increases. Such a relativistic effect further leads to stiffer EOS than the case of LP (1981).
For $\rho_{\rm B}\lesssim\rho_{0}$, the $E$ (two-body) in the present case is less attractive than that of LP (1981) as seen in Fig. 2. The deficit attraction in the $E$ (two-body) is compensated with more attractive TNA than the case of LP (1981). For $\rho_{\rm B}\gtrsim\rho_{0}$, the $E$ (TNA) in both cases decrease in magnitude and become negligible for $\rho_{\rm B}\gtrsim$ 0.4 fm-3.
@@ V.3 Dependence of the EOS for SNM and PNM on the slope $L$
Next one considers dependence of the EOS for SNM on the slope $L$. In Fig. 3, the total energy per baryon as well as each energy contribution in SNM are shown as functions of $\rho_{\rm B}$ by the dotted, solid, and dashed lines for $L$=60, 65, and 70 MeV, respectively. The $E$ (TNR) is commonly taken from the SJM2 t2008 for the cases of $L$=(60, 65, 70) MeV and is represented by the single green solid line.
@@
At high densities beyond $\rho_{0}$, the attraction from the $\sigma$-exchange is saturated, so that the remaining repulsion from the $\omega$-exchange is more remarkable for larger $\omega$-mean field. The increase in repulsive energy from the $\omega$-meson exchange with the increase in $L$ from 60 MeV to 70 MeV, as shown by the arrows for the $E$ (two-body) in Fig. 3, plays a decisive role on stiffening the EOS in SNM. Inversely, as $L$ gets smaller than $L\lesssim 55$ MeV, the repulsive energy from the $\omega$-meson exchange is not large enough to compensate the attractive energy by the $\sigma$-meson exchange beyond $\rho_{0}$, so that the system becomes unbound with the two-body interaction only in the present interaction model. Thus, the strength of the binding energy contribution by the TNA at $\rho_{0}$ and the density profile of the two-body $B$-$B$ interaction via the $\sigma$ and $\omega$ exchanges show subtle correlation through the choice of $L$ in the (MRMF+UTBR+TNA) model.
In contrast, as for the (MRMF+NLSI) model, where many-baryon forces are given by the NLSI terms in place of the UTBR and TNA, the two-body $B$-$B$ interaction via the $\sigma$ and $\omega$ exchanges are determined irrespectively of the choice of $L$, since the total energy does not have isospin-dependent terms like the TNA in the case of the SNM, as is shown in Appendix X.4.
To compare the $L$-dependence of the EOS between the PNM and SNM, the total energy per nucleon, $E$ (total), is shown in Fig. 4 and the pressure $p$ is shown in Fig. 5, as a function of $\rho_{\rm B}$ in the case of $L$=(60, 65, 70) MeV, for both PNM and SNM cases. The shaded region in Fig. 5, corresponding to the region of pressures consistent with the experimental flow data in the case of SNM, is taken from Ref. danielewicz2002 . One can see the EOS gets stiffer for larger $L$ over the densities $\rho_{\rm B}=(1-5)\rho_{0}$ for PNM as well as the case of SNM. Therefore, in the present interaction model, not only the TNR but also the slope $L$ controls the stiffness of the EOS at high densities. This feature will be shown to be commonly applied to hadronic matter with the ($Y$+$K$) phase. (see Sec. VII. )
@@
Figure 5: The pressure in SNM (black curves) and in PNM (green curves) as a function of $\rho_{\rm B}$ for $L$=(60, 65, 70) MeV. The shaded region, corresponding to the region of pressures consistent with the experimental flow data in the case of SNM, is taken from Ref. danielewicz2002 .
In danielewicz2002 , the allowable region in the pressure ($p$)-density ($\rho_{\rm B}$) plane for the EOS in SNM and PNM has been obtained from the analyses of transverse and elliptic flow data in nuclear collisions [The shaded region in Fig. 5 in the case of SNM]. Quantitatively, the result for $L\lesssim$ 65 MeV lie within the upper bound of the experimentally consistent region for $\rho_{\rm B}$ = (2--4) $\rho_{0}$, while those for $L\gtrsim$70 MeV slightly above it. For the latter case, some extra softening effect might be responsible in order to be consistent with the experimental implications.
In Fig. 4, the density-dependence of the symmetry energy, $S(\rho_{\rm B})$, defined by $S(\rho_{\rm B})$=$E$ (total)(PNM)--$E$ (total)(SNM), is shown in the case of $L$=(60, 65, 70) MeV by the red solid lines. $S(\rho_{\rm B})$ increases monotonically with density for a given $L$. Since there is a cancellation of the $L$-dependent part of the total energy between the cases of PNM and SNM, the slope $L$-dependence of $S(\rho_{\rm B})$ is not remarkable even at high densities.
@@ VI Kaon properties in hyperon-mixed matter and onset of KC
Here kaon properties in hyperon-mixed matter are considered, and onset density of KC with the interaction model (ChL+MRMF+UTBR+TNA) is obtained.
@@ VI.1 Meson-hyperon coupling constants for description of hyperon-mixed matter
For the description of hyperon-mixed matter, one sets the values of the meson-hyperon coupling constants to obtain the hyperon-nucleon and hyperon-hyperon interactions in the MRMF.
@@
where $\langle m\rangle_{0}$ ($m$=$\sigma$, $\sigma^{\ast}$, $\omega_{0}$, $R_{0}$, $\phi_{0}$) are the meson mean fields at $\rho_{B}=\rho_{0}$ in SNM, and the last term, which is equal to the second term from the last on the r. h. s. of Eq. (79), comes from the energy density contribution from the UTBR. From recent theoretical and experimental analyses on hypernuclear experiments ghm2016 , the values of the $V_{Y}^{N}$ in Eq. (92) are set to be $V_{\Lambda}^{N}=-27$ MeV, $V_{\Sigma^{-}}^{N}$ = 23.5 MeV, and $V_{\Xi^{-}}^{N}$ = $-14$ MeV, and one obtains $g_{\sigma\Lambda}$, $g_{\sigma\Sigma^{-}}$, and $g_{\sigma\Xi^{-}}$. These coupling constants are obtained for each case of $L$ and listed in Table 2.
@@ VI.2 Onset of $\Lambda$-mixing and composition of matter below the onset density of KC
As is shown below, the $\Lambda$ hyperons always appear at a lower density than KC for various cases of $L$ and $\Sigma_{Kn}$. Therefore first the onset of the $\Lambda$-mixing and composition of matter are considered prior to KC.
@@
Figure 6: (a)  The $\Lambda$ potential $V_{\Lambda}$ (green curves) and ($\mu_{n}-M_{\Lambda}$) (red curves) as functions of baryon number density $\rho_{\rm B}$ for $\Sigma_{Kn}$ = 300 MeV in the case of $L$=60 MeV (dotted lines), 65 MeV (solid lines), and 70 MeV (dashed lines). The filled triangle denotes the onset density of $\Lambda$-mixing, $\rho_{\rm B}^{c}(\Lambda)$, in $\beta$-equilibrated nuclear matter, and the filled circle denotes the onset density of KC, $\rho_{\rm B}^{c}(K^{-})$, in $\beta$-equilibrated hyperon-mixed matter. (b) The same as (a), but for $\Sigma_{Kn}$ = 400 MeV. The profiles for $\Sigma_{Kn}$ = 300 MeV and 400 MeV are the same before appearance of KC. See the text for details.
The onset density of the $\Lambda$ hyperons, $\rho_{\rm B}^{c}(\Lambda)$, in $\beta$-equilibrated neutron-star matter is given by the intersection of the curves for the $\mu_{n}-M_{\Lambda}$ and $V_{\Lambda}$, denoted as the filled triangle. The $\Lambda$ hyperons are mixed over the densities where $\mu_{n}-M_{\Lambda}>V_{\Lambda}$. Numerically $\rho_{\rm B}^{c}(\Lambda)$ is weakly dependent on $L$: $\rho_{\rm B}^{c}(\Lambda)$ = (0.47, 0.43, 0.40) fm-3 for $L$ = (60, 65, 70) MeV.
@@
Table 3: The onset densities at which hyperon-mixing starts and those of KC in the (ChL+MRMF+ UTBR+TNA) model for $\Sigma_{Kn}$ = 300 MeV and 400 MeV in case of $L$=60, 65, and 70 MeV. The $\rho_{\rm B}^{c}(\Lambda)$ is the onset density of $\Lambda$ hyperons in the normal neutron-star matter, $\rho_{\rm B}^{c}(\Xi^{-}\ {\rm in}\ \Lambda)$ the one of $\Xi^{-}$ hyperons in the $\Lambda$-mixed matter, $\rho_{\rm B}^{c}~{}(K^{-})$ the one of KC in the hyperon ($\Lambda$ and/or $\Xi^{-}$)-mixed matter, and $\rho_{\rm B}^{c}(\Xi^{-}\ {\rm in}\ K^{-}\Lambda)$ the one of the $\Xi^{-}$ hyperons in the KC phase in the $\Lambda$-mixed matter. For reference, the central density of the maximum mass star, $\rho_{\rm B,center}(M_{\rm max})$, obtained after solving the T.O.V equation, and the depth of the $K^{-}$ optical potential $U_{K}$, estimated at $\rho_{\rm B}$ = $\rho_{0}$ in the SNM, are listed for each case of $L$ and $\Sigma_{Kn}$.
$L$ $\Sigma_{Kn}$ $U_{K}$ $\rho_{\rm B}^{c}(\Lambda)$ $\rho_{\rm B}^{c}(\Xi^{-}~{}{\rm in}~{}\Lambda)$ $\rho_{\rm B}^{c}(K^{-})$ $\rho_{\rm B}^{c}(\Xi^{-}~{}{\rm in}~{}{K^{-}\Lambda})$ $\rho_{\rm B,center}(M_{\rm max})$
(MeV) (MeV) (MeV) (fm-3) (fm-3) (fm-3) (fm-3) (fm-3)
60 300 --111 0.466 -- 0.598 1.04 1.22
400 --132 0.466 -- 0.486 0.994 1.34
65 300 --111 0.425 0.568 0.653 -- 1.07
400 --131 0.425 -- 0.503 0.900 1.16
70 300 --111 0.397 0.516 0.733 -- 0.966
400 --131 0.397 (0.516) 0.523 0.790 1.01
@@ VI.3 Onset density of kaon condensation in hyperon-mixed matter
The lowest kaon energy $\omega_{K}(\rho_{B})$ decreases as a function of $\rho_{\rm B}$ due to the $K$-$B$ scalar and vector attraction. In the case of the $s$-wave kaon condensation, the $\omega_{K}(\rho_{B})$ intersects with the kaon chemical potential $\mu_{K}$, which is equal to the charge chemical potential $\mu$ in Eq. (74). The onset density $\rho_{B}^{c}(K^{-})$ for the $s$-wave kaon condensation as a continuous phase transition is given by the condition mt92
@@
With the coefficients $d_{p}$ and $d_{n}$, which are determined so as to reproduce the empirical values of the on-shell $s$-wave $K$-$N$ scattering lengths (Appendix X.2), one can see from Eq. (98) that for on-shell $K^{-}$ mesons ($\omega_{K}=m_{K}$), the $s$-wave $K$-$N$ scalar attraction simulated by $\Sigma_{Kb}$ terms is canceled out by each corresponding range term. Also the $\Lambda^{\ast}$-pole, which lies about 30 MeV below the $KN$ threshold, works repulsively to the on-shell $s$-wave $K^{-}$$p$ scattering length. The range terms and $\Lambda^{\ast}$-pole term thus have sizable contributions to the $K^{-}$ self-energy in free space. However, as baryon density increases such that $\rho_{\rm B}\gtrsim\rho_{0}$, the $K^{-}$ energy $\omega_{K}$ decreases due to the $s$-wave $K$-$N$ attractive interactions, getting $\omega_{K}/m_{K}\ll 1$. In such a case, the $\Sigma_{Kb}$ terms and $s$-wave $K$-$N$ vector interaction, $-2X_{0}\omega_{K}$, become dominant over the range terms and the $\Lambda^{\ast}$-pole term [^2] In Ref. fmmt1996 , the same result has been obtained in the second-order perturbation with respect to the axial-vector current (the second-order effect) in the framework of current algebra and PCAC t88 . The present result on the lowest $K^{-}$ energy naturally agrees with those obtained by other works based on the chiral perturbation theory lbm95 ; tw1995 ., and the $K^{-}$ self-energy is well described without the range terms and $\Lambda^{\ast}$ term as
@@
which one uses for actual numerical calculations throughout this paper. The term leading to the $s$-wave $K$-$B$ scalar interaction in the self-energy (99) is proportional to $\rho_{b}^{s}$, which has a complicated $\rho_{\rm B}$-dependence:
@@
A scale of the $s$-wave $K$-$N$ attraction is characterized by the $K^{-}$ optical potential $U_{K}$ in the SNM, which is defined in terms of the $K^{-}$ self-energy [(99)] as $U_{K}=\Pi_{K}(\omega_{K};\rho_{\rm B})/(2\omega_{K})|_{\rho_{\rm B}=\rho_{0}}$. In Table 3, the $U_{K}$ is listed for each case of $L$ and $\Sigma_{Kn}$. The value of $U_{K}$ has sensitive dependence on $\Sigma_{Kn}$, while it depends little on the slope $L$. The deduced value of the depth $|U_{K}|$ (110 MeV--130 MeV) is larger than the theoretical values in the chiral unitary approach ro2000 and the recent optimal value of the real part of the $K^{-}$ optical potential depth $|V_{0}|$ = 80 MeV with the imaginary part $W_{0}$ = --40 MeV, which was deduced from the measured spectrum shape of the inclusive missing-mass spectrum of 12C ($K^{-},p$) reactions in the J-PARC E05 experiment ichikawa2020 . Nevertheless, the result is within the allowable range, considering that a much larger value of $|V_{0}|$ ($\gtrsim$ 150 MeV) is confronted with the difficulty of reproducing the spectrum ichikawa2020 .
@@
Going into details of (i), one shows the density-dependence of the $X_{0}$ [Eq. (16)] and that of the “effective mass” $m_{K}^{\ast}$ of the $K^{-}$ meson [Eq. (65)] in Fig. 7. For reference, the particle fractions $\rho_{a}/\rho_{\rm B}$ and baryon scalar densities $\rho_{b}^{s}$ including the ($Y$+$K$) phase are also shown as functions of $\rho_{\rm B}$ in Fig. 8 and Fig. 9, respectively.
@@
In Refs. panda1995 ; carlson2000 the effect of $K$-$N$ and $N$-$N$ correlations was taken into account in the study of kaons in dense matter, and it was shown that the correlations tend to push up the $\omega_{K}$ over the relevant densities. Subsequently the effect of the short-range correlations was discussed in detail and was shown to be moderate wrw1997 ; ww1997 as compared to the preceding results panda1995 ; carlson2000 . The present result on the $U_{K}$ is similar to that of Refs. wrw1997 ; ww1997 with inclusion of the short-range correlations ($U_{K}\approx-120$MeV), whereas the $\omega_{K}$ in the model is slightly smaller than that of Refs. wrw1997 ; ww1997 .